The density of ammonia gas under certain conditions 0.505 g/L. Calculate its density in g/cm3.
1
Expert's answer
2020-09-14T08:26:01-0400

As 1 L = 1000 cm3, the density of ammonia equals:

0.505 g/L = 0.505 g / 1000 cm3 = 0.000505 g/cm3 = 5.05 × 10-4 g/cm3


Answer: 5.05 × 10-4 g/cm3